OpenStudy (anonymous):

Al gives 1/2 of his marbles to Bev. Bev gives 1/3 of his to Carl. Carl gives 1/4 of these to Dani. If Dani was given 4 marbles, how many did Al have originally?

OpenStudy (lgbasallote):

let x be what al had originally so x/2 was given to bev then bev gives (x/2)/3 = x/6 to carl carl gives (x/6)/4 = x/24 todani it says that x/24 is 4 so x/24 = 4 x = 24 * 4 x = 96 does that help?

OpenStudy (anonymous):

Yes it does thanks, I was trying to figure out how to work backwards but kept getting stuck so thanks for the clarrification

hero (hero):

To post the next question, close this one. Then the "ask a question" box will show up again.

OpenStudy (anonymous):

okay I've got it thanks this is my 1st time using this, well duh, guess u figured that one out already ;) thanks

hero (hero):

Okay, good job

hero (hero):

I don't see where you have posted the next question.

OpenStudy (anonymous):

I did post it not sure why u can't see it because someone asked a question but have not yet replied

hero (hero):

I already responded to the other question.